### v1

# The Power of LLM-Generated Synthetic Data for Stance Detection in Online Political Discussions

Abstract
Stance detection holds great potential for enhancing the quality of online political discussions, as it has shown to be useful for summarizing discussions, detecting misinformation, and evaluating opinion distributions. Usually, transformer-based models are used directly for stance detection, which require large amounts of data. However, the broad range of debate questions in online political discussion creates a variety of possible scenarios that the model is faced with and thus makes data acquisition for model training difficult. In this work, we show how to leverage LLM-generated synthetic data to train and improve stance detection agents for online political discussions: (i) We generate synthetic data for specific debate questions by prompting a Mistral-7B model and show that fine-tuning with the generated synthetic data can substantially improve the performance of stance detection. (ii) We examine the impact of combining synthetic data with the most informative samples from an unlabelled dataset. First, we use the synthetic data to select the most informative samples, second, we combine both these samples and the synthetic data for fine-tuning. This approach reduces labelling effort and consistently surpasses the performance of the baseline model that is trained with fully labeled data. Overall, we show in comprehensive experiments that LLM-generated data greatly improves stance detection performance for online political discussions.
1 Introduction
With the recent advent of powerful generative Large Language Models (LLMs) such as ChatGPT [^1] https://chat.openai.com/, Llama (Touvron et al., 2023) and Mistral (Jiang et al., 2023), investigating how synthetic data from LLMs can be leveraged for downstream tasks has become a very important and relevant topic of research. It is well known that Natural Language Processing (NLP) tasks based on single-label or multi-label classification such as stance and sentiment detection benefit immensely from large amounts of labelled data (Mehrafarin et al., 2022). However, in many cases, labelled data is scarce, especially in the realm of online political discussions: one issue is that gathering labelled data may not be allowed for privacy reasons. Another issue is that especially in stance detection for online political discussions comments and their stances are intrinsically linked to a question. There are a potentially infinite number of specific questions open for debate for which different types of comments (and their stances) arise. This can be a problem, since the samples may not be varied or pertain to only one class, hence usually yielding an unbalanced dataset. For instance, performance drops severely for questions in the test dataset, that are not present in the training dataset (Vamvas and Sennrich, 2020).
An obvious approach to stance detection might be to apply LLMs directly, which has been tried with mixed results (see Section 2). Furthermore, LLMs are not yet fast enough at inference time for smaller real world interactive deployment in online discussions, unless adequate hardware infrastructure is provided. Also, using public hosting infrastructure may not be feasible when use cases require anonymity and privacy.
Positive SynthNegative SynthTrain dataMost informativeRefer to captionRefer to caption AB
Figure 1: We investigate the use of LLM-generated synthetic data for stance detection in online political discussions. (A) We generate synthetic data  for specific questions using a Mistral-7B model. The synthetic data is then used to fine-tune the stance detection model. We show that fine-tuning with synthetic data improves the performance of the model, since the synthetic data is roughly faithful to the real data’s  underlying distribution. However, some real world samples  cannot be captured by the synthetic data. (B) We therefore use the synthetic data to identify the most informative samples  in the unlabelled real data pool. That is, we identify the samples that are on the boundary of the two classes of the synthetic data which are better off labelled by human experts. We show that combining the synthetic data with the manually labelled most informative samples improves the performance of the model even further.
While LLMs might not be directly useful for stance detection, we show in this paper, that they can be used to generate synthetic training data for more specific (and smaller) NLP models, such as BERT (Devlin et al., 2019), to create powerful stance detection for online political discussions that potentially work for any type of question. How exactly synthetic data from an LLM should be combined with a small pool of unlabelled training data will be investigated along the following questions:
Question 1 (Q1): Does fine-tuning with synthetic data improve the stance detection model? Answer: Yes, and the more the better!
One problem of stance detection datasets for online political discussions is the lack of (balanced) data for specific questions (sometimes also called issues). While fine-tuning with synthetic data is possible for BERT-based models, it was previously found to be challenging (Mahmoudi et al., 2024). Using more powerful LLMs, we can show that fine-tuning with synthetic data does improve stance detection performance in online political discussions. This also holds when augmenting existing labelled data.
Question 2 (Q2): Can the labelling effort be reduced by controlling what example to manually label using active learning with synthetic data? Answer: Yes, even if we don’t have real true labels initially.
Another challenge in stance detection for online political discussions is labelling data, which is expensive and time consuming. Active learning can help us to identify the most informative samples that are then labelled manually, even if there are initially no labelled samples. For this we use the synthetic samples which have labels by construction, to select the most ambiguous (and thus most informative) samples in the unlabelled dataset.
Question (Q3): Can fine-tuning using the most informative samples plus the synthetic data outperform fine-tuning with a completely manually labelled dataset? Answer: Yes, it can!
Our study reveals that the most effective way to improve stance detection performance is to combine both synthetic data and the most informative samples obtained via active learning. Surprisingly, this is even better than using a full dataset that has been completely manually labelled.
2 Background
Stance Detection for Online Political Discussions.
Stance detection is a sub-task of sentiment analysis (Romberg and Escher, 2023) and opinion mining (ALDayel and Magdy, 2021) that aims to automatically detect the stance (e.g., favor, against, or neutral) of the author of a contribution towards a discussed issue, or target. In online political discussions, stances can be defined as the side that the author is on, e.g., for or against an increase in value added tax. Research has identified stance detection as an important yet challenging task; it is important because the automated detection of stances can, e.g., improve discussion summarisation (Chowanda et al., 2017), facilitate the detection of misinformation (Hardalov et al., 2022), and provide more comprehensive evaluations of opinion distributions in online political discussions and participation processes (Romberg and Escher, 2023). Stance detection is also implemented as a feature in many recommender systems and discussion platforms (Küçük and Can, 2020). Still, due to its dependency on context, stance detection is a highly challenging task. In online political discussions, for example, detecting the stance of a contribution towards a question requires a proper identification of both the question under discussion and the position of the contributor towards it. However, users regularly deviate from the original question and discuss multiple topics in the same thread (Ziegele et al., 2014), resulting in a high number of potential questions with often very few instances for training purposes. Fine-tuning transformer-based models (Vaswani et al., 2017) to solve stance detection is a common practice, but training these models requires a large amount of annotated data, which for the large variety of questions in online political discussions is unfeasible to acquire. We therefore in our work, show how to improve stance detection for online political discussions with synthetic data.
Active Learning.
The aim of active learning is to minimize the effort of labelling data, while simultaneously maximizing the model’s performance. This is achieved by selecting a query strategy that chooses the most interesting samples from a set of unlabelled data points, which we refer to as most informative samples. These samples are then passed to, e.g., a human annotator for labelling. The usual setup consists of iteratively training a model on a small subset of manually labelled data that stems from a pool of unlabelled data. There exist many different query strategies such as Query By Comittee (QBC, (Seung et al., 1992)), Minimum Expected Entropy (MEE, Holub et al. (2008) or Contrastive Active Learning (CAL, Margatina et al. (2021)). By actively choosing samples and asking for the correct labelling, the model is able to learn from few labelled data points, which is advantageous especially when annotated datasets are not available. Within the domain of political text analysis, many different tasks lack large amounts of annotated data. It has been already shown in the past that these tasks can benefit from the active learning: e.g., stance detection (Kucher et al., 2017), topic modeling (Romberg and Escher, 2022), speech act classification (Schmidt et al., 2023) or toxic comment classification (Miller et al., 2020). In this work, we examine how LLM-generated synthetic data can be used instead of real labelled data to select the most informative samples to be manually labelled.
Using LLM-generated synthetic data for fine tuning.
Recent work has shown that synthetic data generated from LLMs can be used to improve the performance of a model on downstream tasks. Møller et al. (2023) showed that synthetic data can be used to improve the performance of a model on downstream classification tasks by comparing the performance of a model finetuned on LLM-generated data to crowd annotated data. In many cases the model finetuned on LLM-generated data outperforms the model finetuned on crowd annotated data. Mahmoudi et al. (2024) study the use of synthetic data for data augmentation in stance-detection. The authors use GPT-3 to generate synthetic data for a specific topic with mixed results due to the inability of GPT-3 to generate good data samples. In our work,we use a newer LLM model, Mistral-7B, which generates better synthetic data samples and show that we can generate synthetic data that matches the real data distribution. Veselovsky et al. (2023) analyse in which ways synthetic data is best generated for tasks like sarcasm detection and sentiment analysis. The authors reach the conclusion that grounding the prompts to generate the synthetic data to real samples helps improve the quality of the synthetic data. Similarly, Li et al. (2023) argue that subjectivity in the classification task determines whether synthetic data can be used effectively.
Using LLMs directly for stance detection.
It has been shown that LLMs can be used directly for stance detection such as (Cruickshank and Xian Ng, 2023), (Burnham, 2023). However, the general conclusion of these studies is that while LLMs are competitive with other transformer models such as BERT, especially for edge cases, they exhibit replication issues. (Burnham, 2023) also discuss the posibility of pre-training models on more specific data to improve the generalisation capability of the model. We therefore focus on using LLMs to generate synthetic data to solve key challenges in stance detection such as the lack of available data for specific topics and labelling large amounts of data, rather than using LLMs directly for the task.
3 Method
In line with our declared contributions, we present our core ideas to improve the performance of the stance detection model: (i) To fine-tune the model with synthetic data, we first define the synthetic dataset and show our pipeline to generate it. The baseline model is then further fine-tuned on the synthetic data. (ii) We then present our synthetic data-based approach to identify the most informative samples from a batch of unlabelled data. We propose a synthetic extension to the QBC (Query by Commitee, (Seung et al., 1992)) method, where the synthetic data act as an ensemble of experts.
3.1 Preliminaries
Political discussions are typically centered around questions $q\in\mathcal{Q}$ (sometimes also called issues or targets). For stance detection, we usually have for each of these questions $q$ a set of labelled data $\mathcal{D}^{(q)}=\{(x^{(i)},y^{(i)})\}_{i=1}^{I}$ where $x^{(i)}\in\mathcal{X}$ is a statement (or comment) and $y^{(i)}$ is the stance of the statement, with $y^{(i)}\in\{0,1\}=\mathcal{Y}$. Note, that we use the notation $\mathcal{D}^{(q)}$ for labelled and for unlabelled datasets (then the labels are ignored). We view the stance detection model as a binary classification function $f:\mathcal{Q}\times\mathcal{X}\rightarrow\mathcal{Y}$, where we included the question as input to provide context. The classifier $f$ can be split into two parts, $f=h\circ g$: the first part is a pre-trained language model $g:\mathcal{Q}\times\mathcal{X}\rightarrow\mathcal{E}$ such as BERT (Devlin et al., 2019) or RoBERTa (Liu et al., 2019). The second part is a classifier $h:\mathcal{E}\rightarrow\mathcal{Y}$ on top of the pre-trained model $g$. The pre-trained language model $g$ outputs an embedding $e^{(i)}\in\mathcal{E}$ given a question $q$ and an input $x^{(i)}$. The classifier $h$ is usually a linear layer or a multi-layer perceptron, which takes the embedding $e^{(i)}$ from $g$ as input and outputs a prediction $\hat{y}^{(i)}$. The stance detection model is fine-tuned by minimizing the cross-entropy loss between the predicted labels $\hat{y}^{(i)}=f(q,x^{(i)})$ and the actual labels $y^{(i)}$.
3.2 Generating Synthetic Data for Stance Detection
To generate synthetic samples, we employ a quantized version of the Mistral-7B-instruct-v0.1 model [^2] https://huggingface.co/turboderp/Mistral-7B-instruct-exl2 to generate comments on a specific question $q$, using the following prompt:
A user in a discussion forum is debating other users about the following question: [q] The person is in favor about the topic in question. What would the person write? Write from the person’s first person perspective.
where "[q]" must be replaced with the question $q$. Similarly, to generate a negative sample, we replace "is in favor" with "is not in favor". As in the X-Stance dataset (Vamvas and Sennrich, 2020), we assign the two labels 0 and 1. We denote the question-specific synthetic dataset as:
$\mathcal{D}_{\text{synth}}^{(q)}=\big{\{}(x_{\text{synth}}^{(m)},1)\big{\}}_{m=1}^{M/2}\cup\big{\{}(x_{\text{synth}}^{(m)},0)\big{\}}_{m=1+M/2}^{M}$ (1)
where half of the $M$ synthetic data samples have positive labels, i.e., are comments in favor for the posed question, while the other half is against. Since the dataset is in German, we translate the questions $q$ with a "NLLB-300M" (NLLB Team et al., 2022) translation model. The English answers from the Mistral-7B model are then translated back to German using the translation model.
We show the translation pipeline in Figure 9 along with translations of the questions chosen for our experiments in Table 2. We also attach the synthetic generated data in both English and German to the paper. In our experiments, this translation pipeline worked well even for complex topics in online political discussions, making our method applicable to any language where a good translation model exists. Overall, the generated dataset $\mathcal{D}_{\text{synth}}^{(q)}$ will be used in two ways: (i) to augment the existing dataset $\mathcal{D}^{(q)}$ in order to increase the amount of training data, and (ii) to detect the most informative samples in the unlabelled data pool, which is explained next.
3.3 Synthetic Data-driven Query By Comittee (SQBC)
In the typical active learning scenario, we have a set of unlabelled data, where the goal is to choose the most informative samples to label. Our approach is inspired by the two methods Query by Comittee (QBC) (Seung et al., 1992) and Contrastive Active Learning (CAL, Margatina et al. (2021)): instead of using QBC’s ensemble of experts and the KL-divergence based information score in CAL, we directly use the synthetic data and its labels to produce classification scores for the unlabelled data using $k$ nearest neighbors. The most informative samples are then the data points with the most indecisive scores. Since we query a synthetic dataset instead of an ensemble of experts (as in QBC), we call our method Synthetic data-driven QBC or short SQBC.
Our method performs three steps:
(1) Generate the embeddings.
Given some embedding function $g:\mathcal{Q}\times\mathcal{X}\rightarrow\mathcal{E}$, we generate embeddings for the unlabelled data,
$E=\big{\{}e^{(i)}\big{\}}_{i=1}^{I}=\big{\{}g(q,x^{(i)})\big{\}}_{i=1}^{I}$ (2)
and for the labelled synthetic data
$E_{\text{synth}}=\big{\{}e_{\text{synth}}^{(m)}\big{\}}_{m=1}^{M}=\big{\{}g(q,x_{\text{synth}}^{(m)})\big{\}}_{m=1}^{M}.$ (3)
Note that $q$ is the question for which we generate the synthetic data and for which we want to detect the most informative samples. If obvious from the context, we often omit the superscript $(q)$.
(2) Using the synthetic nearest neighbours as oracles to score the unlabelled data.
For the $i$-th unlabelled embedding $e^{(i)}$ let $\text{NN}(i)$ be the set of indices of the $k$ nearest neighbours (wrt. to the embeddings using the cosine similarity) among the labelled embeddings $E_{\text{synth}}$. The score for each unlabelled data point counts the number of labels $y_{\text{synth}}^{(m)}=1$ among the nearest neighbours, i.e.,
$s(i)=\sum_{m\in\text{NN}(i)}y_{\text{synth}}^{(m)}\in\{0,\ldots,k\}.$ (4)
For our experiments, we choose $k=M/2$ which worked well across all experiments (other values for $k$ are possible, but did not lead to significantly better results).
(3) Choosing the most informative samples.
The scores take values between 0 and $k$. For 0, the synthetic nearest neighbours all have labels $y_{\text{synth}}^{(m)}=0$, for value $k$, all have labels $y_{\text{synth}}^{(m)}=1$. The most informative samples have a score around $k/2$. We thus adjust the range of the scores so that values in the middle range have the smallest scores (close to 0). We do this by subtracting $k/2$ from the score and taking the absolute value,
$s^{\prime}(i)=|s(i)-k/2|.$ (5)
The $J$ most informative samples $\mathcal{D}^{(q)}_{\text{MInf}}\subset\mathcal{D}^{(q)}$ among the unlabelled samples are the $J$ samples with the smallest scores. In the experiments we vary $J$ to study the impact of manually labelled most informative samples.
(4) Labelling the most informative samples.
This step must be performed manually, i.e., the most informative samples selected in the previous step are labelled by a human expert. Attempting to assign labels automatically using majority vote from the synthetic samples did not work well, since the most informative samples are exactly at the boundary between the classes, leading to random and thus noisy labels.
4 Experiments
We now describe the datasets, the experiments and the results. Our experiments follow the structure of the questions Q1, Q2 and Q3 posed in the introduction. Apart from presenting numerical results for these questions, we also provide further insights by visualizing the embeddings of the synthetic and real world datasets and by showing the selection behaviour of the active learning methods.
4.1 Datasets
X-Stance dataset.
To evaluate our methods, we use the X-Stance dataset (Vamvas and Sennrich, 2020), a multilingual stance detection dataset, which contains a total of $67,000$ annotated comments on $194$ policy-related questions, answered by Swiss election candidates. X-stance includes comments in German ($48,600$), French ($17,200$) and Italian ($1,400$). For this paper we only included the German comments, but as mentioned above, other languages can be easily included by automatic translation. The data has been extracted from smartvote, a Swiss voting advice platform. The comments are labelled either as being in favor (positive) or against (negative) the posed question, which is a typical labelling scheme for stance detection datasets. The X-Stance dataset is split in training and testing questions, i.e,. a question in the training dataset does not appear in the test dataset. Furthermore, for each question $q$ from the training data, there are several annotated comments, which form the dataset $\mathcal{D}_{\text{train}}^{(q)}$. Similarly, for each question $q$ in the test data, the set of annotated comments is written as $\mathcal{D}_{\text{test}}^{(q)}$. To refer to the whole training dataset we write $\mathcal{D}_{\text{train}}=\cup_{q\in\mathcal{Q}}\mathcal{D}^{(q)}_{\text{train}}$.
To limit computation time, we selected 10 questions from the test dataset to evaluate our method. The selection was done to best reflect the variability of the data (see Appendix D.1): we choose questions with small amount of comments, unbalanced comments and also balanced comments. Furthermore, for 5 of the questions the majority class is favor and for the other 5 the majority class is against.
Synthetic dataset.
To test the synthetic data-augmentation and active learning based on SQBC (see Section 3.3) we generate sets of varying sizes $M=\{200,500,1000\}$ for each of the $10$ questions. The synthetic data follows the same structure as the data from the X-stance dataset, where for a specific question $q$ we have $M$ comments and $M$ labels. Each set contains $M/2$ positive labels and $M/2$ negative labels, i.e., the synthetic data is balanced.
4.2 Experimental Setup
General setup.
For all experiments, the starting model is the pre-trained BERT base [^3] https://huggingface.co/dbmdz/bert-base-german-uncased model. We adapt the model to the stance detection task by training it on the X-Stance training dataset $\mathcal{D}_{\text{train}}$ (across all questions). We call this the Baseline, and it is characteristic for BERT-based stance detection (e.g., Vamvas and Sennrich (2020)). For fine-tuning and testing we evaluate the given model separately on 10 chosen questions from the test dataset of X-Stance for all experiments. For each question $q$ we split $\mathcal{D}_{\text{test}}^{(q)}$ into a $60/40$ train/test split (repeated with 5 different seeds to get error bars) and use the train split for fine-tuning to the given question and the test split for evaluation.
Evaluated methods.
Next, we briefly describe the different evaluated methods. Table 1 shows an overview of which datasets are used for each method/variant:
Baseline: this is the default model only trained on $\mathcal{D}_{\text{train}}$, (e.g., Vamvas and Sennrich (2020)).
   True Labels: this variant assumes that all true labels for the train split of $\mathcal{D}_{\text{test}}^{(q)}$ are available
   and thus performs no active learning. This is the case $\mathcal{D}^{(q)}_{\text{MInf}}=\mathcal{D}^{(q)}_{\text{test}}$ for each $q$ in the test set.
   SQBC: we apply our active learning approach to get the most informative samples $\mathcal{D}_{\text{MInf}}^{(q)}$.
CAL: we apply Contrastive Active Learning (CAL, Margatina et al. (2021)) to get the most
    informative samples $\mathcal{D}_{\text{MInf}}^{(q)}$.
   Random: as an ablation, we randomly select the most informative samples $\mathcal{D}_{\text{MInf}}^{(q)}$.
Note that the "+Synth" variants additionally fine-tune the model on the synthetic data $\mathcal{D}_{\text{synth}}^{(q)}$.
Evaluation.
Our main results report the average F1 score over 10 selected questions from the test dataset evaluated on the comments from the test split. The error bars represent the average standard deviation over the 10 questions for 5 runs with different seeds. More detailed results per question are shown in Appendix E.
4.3 Experiments
For each question proposed in Section 1, we designed a specific experiment to test the impact of LLM-generated synthetic data augmentation for stance detection in online political discussions:
  • Q1: The two models Baseline and True Labels are augmented with synthetic data of varying sizes.
  • Q2: We compare the 3 active learning strategies to determine the most informative samples: CAL, SQBC and Random. To study reduction in labelling effort, we define 4 different budgets of data samples ($75\%,50\%,25\%$ and $10\%$ of $\mathcal{D}^{(q)}_{\text{test}}$) to be selected for manual labelling. These variants are also compared against Baseline and True Labels.
  • Q3: We combine fine-tuning with the synthetic data (Q1) and the (manually labelled) most informative samples (Q2). We apply the same testing procedure as in Q2, where we test the performance for different labelling budgets, while also comparing to Baseline and True Labels.
4.4 Results
Refer to caption Fine-tuning with Synthetic Data
Figure 2: Q1: Fine-tuning the model with synthetic data improves performance for increasing dataset size: Shown are the F1-Score of Baseline (left) and True Labels (right) for increasing synthetic dataset. Even if a dataset has been fully labelled, augmenting it with synthetic data proves equally as effective.
Refer to caption Fine-tuning with most informative samples selected with synthetic data
Figure 3: Q2: SQBC selects the best samples for manual labelling and improves upon the baseline model. Increasing the synthetic data size for the selection of most informative samples only improves performance marginally, which shows that uncertain samples can already be identified with a small synthetic dataset.
Refer to caption Fine-tuning with most informative samples and synthetic data
Figure 4: Q3: Fine-tuning with most informative samples and synthetic data surpasses the model fine-tuned with all true labels: We label less samples manually, while outperforming the model with all true labels. The reason for the performance increase can be attributed on the effects from Q1 and Q2: (i) the synthetic data helps the model learn the underlying distribution. (ii) The most informative samples improve the model where the synthetic data distribution is not expressive enough.
Q1: Fine-tuning with synthetic data reaches on average the performance of the model with all true labels.
Figure 2 shows the effectiveness of fine-tuning with synthetic data by fine-tuning Baseline and True Labels with different amounts of synthetic data $M=\{200,500,1000\}$. The baseline model with $1000$ synthetic data samples almost reaches the average performance of the model with all true labels. Thus, fine-tuning with synthetic data attains similar performance to labelling all data, without having to label any real data sample. Moreover, even if a dataset has been fully labelled, augmenting it with synthetic data improves stance detection performance even further.
Q2: Most informative samples selected by SQBC are more informative than CAL and random selection and do improve upon the baseline model.
In Figure 3, we show that we can reduce labelling effort by using synthetic data to select the most informative samples. In this setting, we only use the synthetic data for active learning and do not fine-tune with the synthetic dataset. On average, SQBC is bettern than CAL and Random. We observe, that fine-tuning only with the most informative samples might not be as effective as fine-tuning with synthetic data, since Baseline+Synth (M=1000) is slightly better than SQBC. However, the advantage of using the most informative samples is that the model is grounded on real data.
Q3: Fine-tuning with the most informative samples and additionally with synthetic data surpasses the model trained with all true labels.
In Figure 4, we show the performance of fine-tuning the baseline model with the manually labelled most informative samples and at the same time with synthetic data. This combines the experiments for Q1 and Q2. We see that on average with as little as $25\%$ of the unlabelled data, we can consistently achieve better performance than the model trained with all true labels (True Labels). For $M=1000$ synthetic samples, we observe that surprisingly random sampling outperforms the other query strategies (Random+Synth). We will discuss this in the next section.
4.5 Interpreting the role of the synthetic data
The synthetic data is at the core of our work. To get a better understanding we analyse the role of synthetic data when using it for data augmentation and for the selection of the most informative samples. To study whether the synthetic data matches the real data, we visualize embeddings of the synthetic data together with the embeddings of the real world data using T-SNE projection (van der Maaten and Hinton, 2008) in Figures 5, and 6. To gain insights into the selection mechanism for different query strategies, we add the T-SNE embeddings of the selected samples in Figures 8 and 7. Again, we discuss our insights along the questions Q1, Q2 and Q3.
Q1: The synthetic data matches the distribution of the real world data.
We visualize the positive and negative synthetic data samples together with the means (cluster centers) of the real world data in Figures 5 and 6. We observe that the synthetic data extends the real world data, which we believe is the reason why fine-tuning with synthetic data is effective in online political discussions. Also, the larger the synthetic dataset size, the more the synthetic data matches the distribution of the real world data. Furthermore, the positive and negative samples are well separated, which we attribute to having pre-trained the BERT-model on $\mathcal{D}_{\text{train}}$ of the X-Stance dataset, providing prior knowledge about the stance detection task.
Q2: SQBC selects samples that are not contained in the synthetic data distribution.
We visualize the selected samples of the SQBC and compare them to the selected samples by CAL and Random in Figures 7 and 8. We see that SQBC selects the unlabelled samples that are mostly in between the two classes of the synthetic data. This is the expected behaviour since we select the samples where the classification score is ambiguous. For random selection, the range of selected samples is broad: some similar samples between the two classes like SQBC are selected, but also within class samples that are not covered by the synthetic data set. This explains why random selection works well with a large synthetic dataset, since it further extends the decision boundary of the model. Finally, CAL selects samples similar to SQBC, but mostly tends to select samples from only one class, resulting in worse performance.
Q3: The synthetic data extends the unlabelled real data, while the most informative samples improve classification by selecting samples not well covered by the synthetic data.
Combining the insights from Q1 and Q2, we see that the synthetic data effectively extends the real world data thus providing a better estimate of the underlying distribution. When adding the most informative samples from the active learning methods, we add samples that are not in the more general synthetic data distribution thus extending the decision boundary of the classifier. We argue that this is why the combination of most informative samples and synthetic samples is so effective.
5 Discussion
Limitations and future work.
One limitation of our approach is the fact that we have to fine-tune a separate model for each question. While this leads to excellent results, a common approach is to fine-tune a single (and thus more general) model for several questions (like Baseline). Having observed that the synthetic data distribution differs (sometimes greatly) for each question, we feel that the extra training is justifiable and necessary. The computing time is not a concern in our case, since fine-tuning the BERT model takes less than a minute even with a synthetic dataset size of $M=1000$ on a reasonable GPU (NVIDIA A100). Future work, could study whether the underlying distribution of the synthetic data can serve as a proxy to combine real unlabelled data from different questions with similar (compatible) underlying distributions. This might enable us to train capable general stance detection models.
Conclusion.
In this work, we presented how to improve stance detection models for online political discussions utilizing LLM-generated synthetic data: (i) we showed that fine-tuning with synthetic data related to the question improves the performance of the stance detection model. After analysing the embeddings of the synthetic data, we attribute this to the LLM-generated synthetic data aligning well with the real data for the given question. (ii) We reduce the labelling effort by selecting the most informative samples informed by synthetic data with SQBC. The synthetic data are separated by construction into the positive and negative classes, which makes finding the most informative samples more straightforward, since we can count the labels of the k-nearest synthetic neighbours to determine classification ambiguity. (iii) Fine-tuning with synthetic data can be further enhanced by adding the most informative samples, outperforming the method that uses all true labels. Concluding, our findings demonstrate the potential of synthetic data to significantly improve the efficiency and effectiveness of stance detection in online political discussions.
Appendix A Broader Impact
Our work has the potential to improve the efficiency and effectiveness of stance detection in online political discussions. This can be beneficial for social media platforms, news outlets, and political campaigns, where the detection of stance can help to understand the sentiment of the users. The use of synthetic data in this case is not critical for the user, since the synthetic data is only used to improve the model’s performance. Stance detection could also be used to for bad purposes, e.g., to manipulate the public opinion. However, our work does not contribute to this, since we only focus on improving the model’s performance, i.e., the model’s ability to detect stance.
Appendix B Visualizations
B.1 Visualizing the synthetic data
We visualize the synthetic data together with the real world data for $M=1000$ and $M=200$ in Figures 5 and 6. We plot the data points of the synthetic data in blue and red for the positive and negative samples, respectively. The means of the real world data are plotted as a regular polygon with 8 sides. We observe that the synthetic data extends the real world data, which we consider a factor as to why fine-tuning with synthetic data is effective in online political discussions. Also, the larger the synthetic dataset size, the more the synthetic data matches the distribution of the real world data since for $M=200$ (see Figure 6) the mean are not as well aligned with the synthetic data. Furthermore, the positive and negative samples are well separated, which we attribute to having pre-trained the BERT-model on $\mathcal{D}_{\text{train}}$ of the X-Stance dataset, giving the prior knowledge about the stance detection task.
Question 1Refer to caption
Question 2Refer to caption
Question 3Refer to caption
Question 4Refer to caption
Question 5Refer to caption
Question 6Refer to caption
Question 7Refer to caption
Question 8Refer to caption
Question 9Refer to caption
Negative SyntheticPositive SyntheticPositive Train(Mean)Negative Train(Mean)
Figure 5: Visualization of synthetic data with train data means for $M=1000$ synthetic data. For a larger synthetic dataset size, the means of the synthetic data are well aligned with the real world data and the positive and negative samples are well separated. The synthetic data thus extends the real world data, which we consider a factor as to why fine-tuning with synthetic data is effective in online political discussions.
Question 1Refer to caption
Question 2Refer to caption
Question 3Refer to caption
Question 4Refer to caption
Question 5Refer to caption
Question 6Refer to caption
Question 7Refer to caption
Question 8Refer to caption
Question 9Refer to caption
Negative SyntheticPositive SyntheticPositive Train(Mean)Negative Train(Mean)
Figure 6: Visualization of synthetic data with train data means for $M=200$ synthetic data: For a smaller synthetic dataset size, the means of the synthetic data are not as well aligned with the real world data as for $M=1000$. However, the positive and negative samples are still well separated.
B.2 Visualizing the query strategies of the active learning methods
We visualize the selected samples of SQBC, CAL and Random query strategies for $M=1000$ and $M=200$ in Figures 7 and 8. We plot the selected samples of the unlabelled data in green. The positive and negative synthetic data samples are plotted in blue and red, respectively. The selected samples are highlighted in orange. We observe that SQBC selects the unlabelled samples that are mostly in between the two classes of the synthetic data. This is the expected behaviour since we select the samples where the classification score is ambiguous. For Random, the range of selected samples is broad: some similar samples between the two classes like SQBC are selected, but also within class samples that are not covered by the synthetic data set. This explains why random selection works well with a large synthetic dataset, since it further extends the decision boundary of the model. For the smaller synthetic dataset $M=200$, the random selection is not as effective, since the selected samples are spread out over the whole data space and not necessarily in between the two classes as with the larger synthetic dataset. Finally, CAL selects samples similar to SQBC, but mostly tends to select samples from only one class.
SQBCRandom SamplingCALPositive SyntheticNegative SyntheticTrain dataSelected
Figure 7: Visualization of SQBC, Random and CAL query strategies for $M=1000$ synthetic data: SQBC selects the unlabelled samples that are mostly in between the two classes of the synthetic data. This is the expected behaviour since we select the samples where the classification score is ambiguous. For random selection, the range of selected samples is broad: some similar samples between the two classes like SQBC are selected, but also within class samples that are not covered by the synthetic data set. This explains why random selection works well with a large synthetic dataset, since it further extends the decision boundary of the model. Finally, CAL selects samples similar to SQBC, but mostly tends to select samples from only one class, resulting in worse performance.
SQBCRandom SamplingCALPositive SyntheticNegative SyntheticTrain dataSelected
Figure 8: Visualization of SQBC, Random and CAL query strategies for $M=200$ synthetic data: For a smaller synthetic dataset size, SQBC is still able to select the unlabelled samples that are mostly in between the two classes of the synthetic data. For Random we see that the selected samples are a bit further away from the synthetic data distribution, which is why we argue it does not perform as well as with the larger synthetic dataset.
Appendix C Additional Experimental Details
C.1 Compute and Runtime
We conduct our experiments on a single NVIDIA A100 80GB GPU and a 32 core CPU. With this setup, for Mistral-7B, the generation of synthetic data takes approximately $3$ hours per question for a synthetic dataset size of $M=1000$. Fine-tuning the BERT model with the synthetic data takes less than a minute. For the active learning methods, the selection of the most informative samples takes less than a minute. Hence the largest computational effort is the generation of the synthetic data.
C.2 Translation of the X-Stance dataset for synthetic data generation
In Figure 9 we show the pipeline for translating the X-Stance dataset for synthetic data generation. We start with a question $q$ from the X-Stance test dataset and translate the question to English with a NLLB-330M model [NLLB Team et al., 2022]. Then we let the Mistral-7B model generate synthetic data, i.e., comments for the translated question. The generated comments are then translated back to German to be used for fine-tuning the model in our experiments.
C.3 Overview of used datasets
In Table 1 we show an overview of the datasets used in our experiments for the different methods we evaluate.
Figure 9: Overview of the pipeline for active learning with synthetic data: We start with a question $q$ from the X-Stance test dataset and translate the question to English with a NLLB-330M model [NLLB Team et al., 2022]. Then we let the Mistral-7B model generate synthetic data, i.e., comments for the translated question. The generated comments are then translated back to German to be used for fine-tuning the model in our experiments.
Config Datasets Manual labels $D_{\text{MInf}}$ True Labels $\mathcal{D}_{\text{t}}$ Synth Aug $D_{\text{synth}}$
Baseline
Baseline + Synth
Oracle
Oracle + Synth
SQBC
SQBC + Synth
CAL
CAL + Synth
Random ✓(randomly selected)
Random + Synth ✓(randomly selected)
Table 1: Synth: Synthetic Data, Aug: Augmentation. We compare different variants of active learning with synthetic data.
Appendix D Dataset
X-Stance is a multilingual stance detection dataset, including comments in German ($48,600$), French ($17,200$) and Italian ($1,400$) on political questions, answered by election candidates from Switzerland. The data has been extracted from smartvote [^4] https://www.smartvote.ch/, a Swiss voting advice platform. For the task of cross-topic stance detection the data is split into a training set, including questions on 10 political topics, and a test set with questions on two topics that have been held out, namely healthcare and political system.
D.1 Chosen questions and their distribution
We present the $10$ chosen questions for our experiments in Table 2. We show the original questions in German and their corresponding English translations by the translation model. Furthermore, we also show the (60 / 40 ) train/test split for each question in Figure 10. We chose 10 questions that reflect the overall distribution of $\mathcal{D}^{(q)}_{\text{test}}$. We chose unbalanced, balanced and low sample size questions to evaluate the effectiveness of our approach.
Figure 10: Distribution of the positive and negative samples for the train and test split of $\mathcal{D}^{(q)}_{\text{test}}$: We show the distribution of the positive and negative samples of the X-Stance test dataset for the questions Q1-Q10. We also show the 60/40 train/test split for the 10 questions. We chose 10 questions that reflect the overall distribution of $\mathcal{D}^{(q)}_{\text{test}}$. We chose unbalanced, balanced and low sample size questions to evaluate the effectiveness of our approach.
Question in German Question in English
Sollen sich die Versicherten stärker an den Gesundheitskosten beteiligen (z.B. Erhöhung der Mindestfranchise) Should insured persons contribute more to health costs (e.g. increase in the minimum deductible)?
Befürworten Sie ein generelles Werbeverbot für Alkohol und Tabak? Do you support a general ban on advertising alcohol and tobacco?
Soll eine Impfpflicht für Kinder gemäss dem schweizerischen Impfplan eingeführt werden? Should compulsory vaccination of children be introduced in accordance with the Swiss vaccination schedule?
Soll die Aufenthaltserlaubnis für Migrant/innen aus Nicht-EU/EFTA-Staaten schweizweit an die Erfüllung verbindlicher Integrationsvereinbarungen geknüpft werden? Should the residence permit for migrants from non-EU/EFTA countries be linked to the fulfilment of binding integration agreements throughout Switzerland?
Soll der Bund erneuerbare Energien stärker fördern? Should the federal government promote renewable energy more?
Befürworten Sie eine strengere Kontrolle der Lohngleichheit von Frauen und Männern? Do you support stricter control of gender pay?
Soll der Bund die finanzielle Unterstützung für die berufliche Weiterbildung und Umschulung ausbauen? Should the Federal Government expand its financial support for continuing vocational training and retraining?
Sollen in der Schweiz vermehrt Spitäler geschlossen werden, um die Kosten im Gesundheitsbereich zu senken? Should more hospitals be closed in Switzerland to reduce healthcare costs?
Eine eidenössische Volksinitiative verlangt, dass der Bundesrat direkt vom Volk gewählt werden soll. Unterstützen Sie dieses Anliegen? A sworn federal people’s initiative demands that the Federal Council be elected directly by the people. Do you support this concern?
Soll die Einführung der elektronischen Stimmabgabe bei Wahlen und Abstimmungen (E-Voting) weiter vorangetrieben werden? Should the introduction of electronic voting in elections and voting (e-voting) be further promoted?
Table 2: Chosen questions for stance detection in German and their English translation
Appendix E Extended Results
We present the extended results for the different synthetic dataset sizes $M=200$, $M=500$ and $M=1000$ in Figures 11, 12 and 13. As in Figure 4, we show the results for the different active learning methods and the different configurations of the synthetic data, while varying the amount of samples that need to be labelled. We compare all methods to True Labels, hence the horizontal line corresponds to the performance of the baseline model fine-tuned with the true labels.
Figure 11: Extended results of Figure 4 for M=200:
Figure 12: Extended results of Figure 4 for M=500
Figure 13: Extended results of Figure 4 for M=1000